OpenStudy (anonymous):

A floor has two square-shaped designs. The area of the second square-shaped design is sixteen times greater than the area of the first square-shaped design. Which statement gives the correct relationship between the lengths of the sides of the two squares?

OpenStudy (anonymous):

if A and B are the areas of the two squares with A being the bigger square And, a^2 = A b^2 = B Then, A = 16B => sqrt(A) = sqrt(16B) a = 4b Hope that answers your question!

OpenStudy (anonymous):

I think I got it. Would the answer be, "The length of the side of the second square is 8 times greater than the length of the side of the first square."?

OpenStudy (anonymous):

Am! That would be 4 times, I think. That's why when we calculate the area (since we square the length of one side), 4 times 4 becomes 16. to if length1 = 4*length2 that implies that: area1 = 16*area2

OpenStudy (anonymous):

two floors consider a & b now b=16a area of square = side square so side b = sqrt b and side a=sqrt 16a =4 sqrt a so 4a =b